Evaluate the limit as x turns to 0

Lim [√(5+x) - √5 / x]


1
Expert's answer
2021-09-06T19:18:50-0400

"lim_{x\to 0}\frac{\sqrt{5+x}-\sqrt{5}}{x}=lim_{x\to 0}\frac{(\sqrt{5+x}-\sqrt{5})(\sqrt{5+x}+\sqrt{5})}{x(\sqrt{5+x}+\sqrt{5})}="

"=lim_{x\to 0}\frac{x}{x(\sqrt{5+x}+\sqrt{5})}=lim_{x\to 0}\frac{1}{\sqrt{5+x}+\sqrt{5}}=\frac{1}{2\sqrt{5}}."